shirley is presently 2 years more than 4 times roses age in 12 years rose will be 1/2 of shirleys age how old are they presently?
Found 2 solutions by lwsshak3, macston:
Answer by lwsshak3(11628)

shirley is presently 2 years more than 4 times roses age in 12 years rose will be 1/2 of shirleys age how old are they presently?
***
let x=Rose's age
4x+2=Shirley's age
..
in 12 yrs:
x+12=Rose's age
4x+2+12=4x+14=Shirley's age
..
x+12=1/2(4x+14)
x+12=2x+7
x=5
4x+2=22
Rose's age=5
Shirley's age=22

Answer by macston(5194)

r=Rose's age; s=Shirley's age=4r+2 years
r+12=1/2(s+12) Substitute for s.
r+12=1/2(4r+2+12) Multiply each side by 2.
2r+24=4r+14 subtract 14 from each side.
2r+10=4r Subtract 2r from each side.
10 years=2r Divide each side by 2
5 years=r ANSWER 1: Rose is 5 years old.
s=4r+2=4(5)+2=22 years ANSWER 2: Shirley is 22 years old.
CHECK:
r+12=1/2(s+12)
5+12=1/2(22+12)
17=1/2(34)
17=17